OpenStudy (anonymous):

What is the sum of a 23-term arithmetic sequence where the first term is 9 and the last term is 97? help ASAP

OpenStudy (ash2326):

It's given \[a_1=9\ \text{( First term)}\] \[a_{23}=97\ \text{( Last term)}\] Now Sum of the series Sn is given by the following relation \[S_n= \frac{n}{2} \times ( a_1+a_{23})\] n=23 Can you solve now?